### v2

# Oblique photons, plasmons, and current-plasmons in relativistic plasmas and their topological implications

Abstract
Photons in vacuum are transverse in any inertial frame; longitudinal photons only exist virtually. By developing a manifestly covariant theory for electromagnetic excitations in relativistic plasmas and applying Wigner’s little group method for elementary particle classifications, we show that photons in plasmas are neither transverse nor longitudinal; they are oblique. Plasmons are electromagnetic and oblique as well. The Lorentz invariant characteristics that distinguishes photons and plasmons is covariant compressibility. The manifestly covariant theory predicts the existence of the current-plasmon, a third oblique, electromagnetic eigenmode, and it also enables the study of photon topology in plasmas. Plasmas remove the photon’s Dirac point in vacuum by giving it an effective mass, but create a tilted Dirac-Weyl point by reviving the virtual longitudinal photon. The manifest covariance of the theory demonstrates that relativistic transparency, despite being widely studied, does not exist in plasmas.
Photons in vacuum are transverse in any inertial frame; longitudinal photons only exist virtually. But we show that in plasmas, media with freely moving charged particles, the characterization of transversality and longitudinality for the electromagnetic excitations are not covariant. In plasmas, photons are neither transverse nor longitudinal; they are oblique. Plasmons, which correspond to the virtual longitudinal photons in vacuum, are oblique and electromagnetic as well.
Only in the special inertial frame co-moving with the plasma, photons are transverse and plasmons are longitudinal and electrostatic. However, albeit non-covariant, this is how the classical theory of photons and plasmons in plasmas is developed. The shortcomings of this classical, non-covariant approach are well-known. For example, the plasmon, a.k.a. the electrostatic Langmuir wave, will carry a time-dependent magnetic component in a general inertial frame, where the electrostatic assumption in the non-covariant theory breaks down. Nevertheless, for an electrostatic plasmon observed in a rest frame, it is still important to know what kind of electromagnetic excitation it would be and whether it is a well-defined eigenmode in a moving inertial frame.
To address these questions, we develop a systematic, manifestly covariant theory for electromagnetic excitations in plasmas. To focus on the main interests of the present study, the photons and plasmons, we take the minimal theoretical model—an unmagnetized, homogeneous, cold plasma. For the theory to be covariant, the plasma carries a general 4-velocity $u_{0}=\left(\gamma,\gamma\beta_{1},\gamma\beta_{2},\gamma\beta_{3}\right)^{T}$ for all charged species, and no a priori assumptions about the polarizations of the modes are allowed. For instance, we should not assume any mode is transverse, longitudinal, or electrostatic. In fact, as we will show, none of the covariant modes are. We don’t even know a priori if the covariant theory admits eigenmodes that can be identified as photons or plasmons as described by the standard non-covariant theory.
Previous work has considered the electromagnetic waves and electrostatic waves in relativistic plasmas [1, 2, 3, 4, 5, 6]. However, despite our best effort, we have found no systematic, manifestly covariant treatment of the subject in the existing literature. Most previous studies focused on relativistic corrections or effects on specific modes, and dispersion relations of the modes were often written down by heuristic arguments based on the requirement of covariance or the necessity of the relativistic factor $\gamma$ and with a priori assumptions about the dispersion relations and polarization properties. Because of the lacking of a systematic, manifestly covariant treatment, certain important phenomena, such as the self-induced transparency in relativistic plasmas, have been debated about in the literature [1, 2, 7, 8, 3, 4, 5, 9, 6]. With the arrival of increasingly powerful laser facilities, we hope the results reported in the present study contribute to the understanding on this subject.
The main obstacle to overcome for a manifestly covariant theory is how to model the collective electromagnetic excitations in plasmas using covariant fields. The standard governing equations expressed in terms of the electromagnetic fields $\boldsymbol{E}$ and $\boldsymbol{B}$, plasma velocity $\boldsymbol{v}$, and plasma density $n$ are not manifestly covariant, because $\boldsymbol{E}$, $\boldsymbol{B}$, and $\boldsymbol{v}$ are 3D vectors instead of covariant vectors or forms in the 4D spacetime. For a manifestly covariant theory, it is necessary to use exclusively the 4-potentional 1-form $A$ or the electromagnetic tensor $F\equiv dA$ and the 4-velocity $u$. The difficulties associated with using $A$ and $u$ are two-fold. The unfixed gauge freedom may introduce nonphysical modes that need to be discarded, and the on-shell condition of $u$ imposes a constraint on allowed eigenmodes in the classical regime. For the present study, it happens that the nonphysical gauge modes and off-shell virtual modes can be easily identified and removed.
Another benefit of the manifestly covariant theory is that symmetry group methods can be applied. We will apply Wigner’s little group method for the classification of elementary particles to classify the electromagnetic excitations in plasmas.
Using the manifestly covariant theory and classification techniques developed, we rigorously show that the system admits one 2D mode that can be identified as a photon and one 1D mode that can be identified as a plasmon, as well as their anti-particles. But the photons are neither transverse nor longitudinal. The electric field of photons is oblique to its momentum. Plasmons are not electrostatic, and they are also oblique. In addition, the manifestly covariant theory also predicts the existence of the current-plasmon, which is a third, previously unknown oblique electromagnetic excitation in relativistic plasmas. If photons, plasmons, and current-plasmons all are electromagnetic and oblique in general inertial frames, is there a covariant label that can be used to distinguish them? We find that covariant compressibility $\chi$ is such a label. Photons and current-plasmons are not covariantly compressible, but plasmons are. The manifestly covariant theory also enables the study of photon topology in plasmas. We will show that plasmas remove the photon’s upright Dirac point in vacuum by giving it an effective mass, but create a tilted Dirac-Weyl point by reviving the virtual longitudinal photon. In addition, the manifest covariance of the theory demonstrates that relativistic transparency [1, 2, 7, 8, 3, 4, 5, 9, 6], despite being widely studied, does not exist in plasmas.
Our starting point is the following manifestly covariant system governing the electromagnetic excitations in an unmagnetized cold plasma,
$\partial_{\mu}J_{s}^{\mu}$ $=0,$ (1)
$\partial_{\mu}F^{\mu\nu}$ $=-4\pi\sum_{s}J_{s}^{\nu},$ (2)
$u_{s\mu}F^{\mu\nu}$ $=-\frac{m_{s}c^{2}}{q_{s}}u_{s\mu}\partial^{\mu}u_{s}^{\nu},$ (3)
where subscript “$s$” denotes particle species, $q_{s}$ and $m_{s}$ are the charge and mass of species $s$ in the rest frame, $u_{s}^{\mu}=\left(\gamma_{s},\gamma_{s}v_{s}^{i}/c\right)$ is the 4-velocity of the species $s$, $n_{s}$ is its density in the momentarily co-moving reference frame (MCRF) [10], and $J_{s}^{\mu}\equiv q_{s}n_{s}u_{s}^{\mu}$ is the 4-current. The sign and index conventions are the same as Misner, Thorn, and Wheeler [11].
The electromagnetic tensor $F$ is defined by the 1-form potential as $F\equiv dA.$ We will take the Lorenz gauge $\partial_{\mu}A^{\mu}=0$, which is Lorentz invariant, i.e., covariant. In terms of $A,$ Eqs. (2) and (3) are, respectively,
$\partial_{\mu}\partial^{\mu}A^{\nu}$ $=-4\pi\sum_{s}J_{s}^{\nu},$ (4)
$2u_{s\mu}\partial^{[\mu}A^{\nu]}$ $=-\frac{m_{s}c^{2}}{q_{s}}u_{s\mu}\partial^{\mu}u_{s}^{\nu}.$ (5)
The unperturbed system is assumed to be uniform over spacetime and has a vanishing 4-potential, i.e., $A_{0}=0.$ Each charged species has a constant equilibrium 4-velocity $u_{s0}$ and a constant equilibrium density $n_{s0}.$ The unperturbed system does not carry net 4-current. To simplify the problem to the bare minimum for our present interest, it is assumed that the equilibrium 4-velocity is the same for all species,
$u_{s0}=u_{0}=\left(\gamma,\gamma\beta_{1},\gamma\beta_{2},\gamma\beta_{3}\right)^{T}.$
The electromagnetic excitation is considered to be a small perturbation of the system. We further assume that the ions are heavy and only electrons are perturbed, i.e.,
$n_{e}$ $=n_{e0}+\delta n,$ (6)
$u_{e}$ $=u_{0}+\delta u,$ (7)
$A$ $=\delta A.$ (8)
The linearized system governing the dynamics of $\delta n$, $\delta u$, and $\delta A$ is
$u_{0}^{\mu}\partial_{\mu}\delta n+n_{0}\partial_{\mu}\delta u^{\mu}$ $=0,$ (9)
$2u_{0\mu}\partial^{[\mu}\delta A^{\nu]}+\frac{m_{e}c^{2}}{q_{e}}u_{0\mu}\partial^{\mu}\delta u^{\nu}$ $=0,$ (10)
$\partial_{\mu}\partial^{\mu}\delta A^{\nu}+4\pi q_{e}\left(\delta nu_{0}^{\nu}+n_{e0}\delta u^{\nu}\right)$ $=0.$ (11)
Let $\delta A^{\mu},$ $\delta u^{\mu},$ $\delta n$ $\sim\exp\left(ik_{\mu}x^{\mu}\right),$ where $k_{\mu}=(-\omega,k_{1},k_{2},k_{3}).$ The system reduces to
$M\left(k,u_{0},\omega_{p}\right)\psi=0,$ (12)
$\psi\equiv\left(\delta A^{\mu},\delta u^{\mu},\delta n\right)^{T},$ (13)
$M\left(k,u_{0},\omega_{p}\right)\equiv\left(\begin{array}[]{ccc}-k_{\mu}k^{\mu}I_{4\times 4}&\omega_{p}^{2}I_{4\times 4}&\omega_{p}^{2}u_{0}^{\mu}\\ u_{0\mu}k^{\mu}I_{4\times 4}-k^{\mu}u_{0\nu}&u_{0\mu}k^{\mu}I_{4\times 4}&0\\ 0&k_{\mu}&u_{0}^{\mu}k_{\mu}\end{array}\right),$ (17)
where $M\left(k,u_{0},\omega_{p}\right)$ is a $9\times 9$ matrix and $\psi$ is the perturbed fields $\delta A^{\mu},$ $\delta u^{\mu},$ and $\delta n$ arranged into a $9\times 1$ column vector. We will also denote $\psi$ by
$\left\{\left(\delta A^{0},\delta A^{1},\delta A^{2},\delta A^{3}\right),\left(\delta u^{0},\delta u^{1},\delta u^{2},\delta u^{3}\right),\delta n\right\}$ (18)
for easy identification of its components. The superscript “$\mu$” in $\delta A^{\mu}$ and $\delta u^{\mu}$ indicates that they are the contravariant components of the perturbed field. In addition, $\delta A$ has been normalized by $m_{e}c^{2}/q_{e}$ and $\delta n$ by $n_{e0}$, and $\omega_{p}^{2}\equiv 4\pi n_{e0}q_{e}^{2}/m_{e}c^{3}$ is the normalized plasma frequency-squared. Note that $n_{e0}$ and $m_{e}$ are quantities measured in the MCRF. The dispersion relation of the eigenmodes is thus
$D\left(k,u_{0},\omega_{p}\right)\equiv\det\left(M\left(k,u_{0},\omega_{p}\right)\right)=0.$ (19)
For a given $\left(u_{0},\omega_{p}\right),$ each root of $D\left(k,u_{0},\omega_{p}\right)$ is an eigen 4-wavevector of the system, and the kernel (null space) of $M$ at each eigen 4-wavevector $k$ is the corresponding eigenvector space. Note that not all vectors in the kernel of $M$ at a given eigen 4-wavevector $k$ are physically valid eigenvectors. Two constraints need to be satisfied, the Lorenz gauge condition,
$k_{\mu}\delta A^{\mu}=0,$ (20)
and the linearized on-shell condition (normalization condition),
$u_{0\mu}\delta u^{\mu}=0.$ (21)
The eigen 4-wavevectors $k$ and eigenvectors $\psi$ of the system determined by Eq. (12) are algebraically complicated. Direct calculations are formidable. However, the system can be significantly simplified by choosing a specific inertial frame. This is enabled by the fact that the system is manifestly covariant, which can described as follows. For a Lorentz transformation
$x^{\nu}\rightarrow x^{\prime\mu}=\Lambda_{\ \nu}^{\mu}x^{\nu},$ (22)
specified by a Lorentz matrix $\Lambda_{\ \nu}^{\mu},$ the 4-vector fields of $u_{0},$ $k,$ $\delta u$, and $\delta A$ in the original frame $x$ transform into $u_{0}^{\prime},$ $k^{\prime},$ $\delta u^{\prime}$, and $\delta A^{\prime}$ as contravariant 4-vectors in the frame $x^{\prime}.$ The density field $\delta n$ is a scalar because it is always measured in the MCRF. It is straightforward to verify that
$\Lambda_{9}$ $M\left(k,u_{0},\omega_{p}\right)\psi=M\left(k^{\prime},u_{0}^{\prime},\omega_{p}\right)\psi^{\prime},$ (23)
$\psi^{\prime}$ $\equiv\Lambda_{9}\psi,$ (24)
$\Lambda_{9}$ $\equiv\left(\begin{array}[]{ccc}\Lambda_{\ \nu}^{\mu}\\ &\Lambda_{\ \nu}^{\mu}\\ &&1\end{array}\right).$ (28)
We will also use the following shorthand notations,
$M$ $\equiv M\left(k,u_{0},\omega_{p}\right),\thinspace\thinspace\thinspace M^{\prime}\equiv M\left(k^{\prime},u_{0}^{\prime},\omega_{p}\right),$ (29)
$D$ $\equiv\det\left(M\right),\thinspace\thinspace\thinspace D^{\prime}\equiv\det\left(M^{\prime}\right).$ (30)
Equation (23) implies that $M$ and $M^{\prime}$ are similar, i.e.,
$\Lambda_{9}M=M^{\prime}\Lambda_{9},$ (31)
and thus
$D=D^{\prime}.$ (32)
For each eigen 4-wavevector $k$ of $D,$ $k^{\prime}\equiv\Lambda k$ is an eigen 4-wavevector of $D^{\prime},$ and for each vector $\psi$ in the kernel of $M$ at $k,$ $\psi^{\prime}\equiv\Lambda_{9}\psi$ is a vector in the kernel of $M^{\prime}$ at $k^{\prime}$. The isomorphism between $\left(k^{\prime},\psi^{\prime},M^{\prime}\right)$ and $\left(k,\psi,M\right)$ reflects the fact that eigenmodes $\left(k^{\prime},\psi^{\prime}\right)$ and $\left(k,\psi\right)$ are actually one identical physical mode seen by two different observers in the frame $x^{\prime}$ and the frame $x$, respectively. With this isomorphism, we can label the physical modes in the frame $x$ by their representation in a specially chosen frame $x^{\prime}$ where the eigen 4-wavevectors $k^{\prime}$ are easier to solve for. It is worthwhile to emphasize that such an isomorphism is not only for the purpose of calculation and classification of eigenmodes, but also a necessity required by the covariance. Eigenmodes in different frames linked by the isomorphism must be treated as one identical mode according to the laws of relativity. As often is the case, preserving the symmetry required by the physics simplifies the seemingly complex calculations.
It is evident that in a frame boosted with a 4-velocity equal to the equilibrium 4-velocity $u_{0}$, the equilibrium plasma will be at rest, i.e.,
$u_{0}^{\prime}=(1,0,0,0)^{T}.$ (33)
The Lorentz matrix for such a boost is
$\Lambda\left(\beta_{1},\beta_{2},\beta_{3}\right)=\left(\begin{array}[]{cccc}\gamma&-\gamma\beta_{1}&-\gamma\beta_{2}&-\gamma\beta_{3}\\ -\gamma\beta_{1}&1+\left(\gamma-1\right)\frac{\beta_{1}^{2}}{\beta^{2}}&\left(\gamma-1\right)\frac{\beta_{1}\beta_{2}}{\beta^{2}}&\left(\gamma-1\right)\frac{\beta_{1}\beta_{3}}{\beta^{2}}\\ -\gamma\beta_{2}&\left(\gamma-1\right)\frac{\beta_{1}\beta_{2}}{\beta^{2}}&1+\left(\gamma-1\right)\frac{\beta_{2}^{2}}{\beta^{2}}&\left(\gamma-1\right)\frac{\beta_{2}\beta_{3}}{\beta^{2}}\\ -\gamma\beta_{3}&\left(\gamma-1\right)\frac{\beta_{1}\beta_{3}}{\beta^{2}}&\left(\gamma-1\right)\frac{\beta_{2}\beta_{3}}{\beta^{2}}&1+\left(\gamma-1\right)\frac{\beta_{3}^{2}}{\beta^{2}}\end{array}\right),$ (34)
where $\beta^{2}\equiv\sum_{j}\beta_{j}^{2}$. We will also refer to the original frame $x$ as the lab frame, where the plasma has a nontrivial equilibrium 4-velocity $u_{0}$. The frame $x^{\prime}$ will be referred to as a rest frame for the obvious reason.
We now further transform the inertial frame $x^{\prime}$ by an additional spatial rotation, i.e.,
$x^{\prime\nu}\rightarrow x^{\prime\prime\mu}=\Lambda_{\ \nu}^{\prime\mu}x^{\prime\nu},$ (35)
where $\Lambda^{\prime}\in\text{SO}(3)\subset\text{SO}^{+}(1,3)$ does not change the 0-th component of $x^{\prime}.$ We choose $\Lambda^{\prime}$ such that in the frame $x^{\prime\prime}$, the spatial component of $k^{\prime\prime}$ is in the 1st-direction,
$k^{\prime\prime}=\left(\Omega,K,0,0\right)^{T}.$ (36)
Here, we use $\Omega$ and $K$ to denote the mode frequency and wavenumber in the frame $x^{\prime\prime}.$ Because $\Lambda^{\prime}\in\text{SO}(3)$, we have
$u_{0}^{\prime\prime}$ $=u_{0}^{\prime}=\left(1,0,0,0\right)^{T},$ (37)
$\Omega$ $=\omega^{\prime}=\Lambda_{\ \mu}^{0}k^{\mu},$ (38)
$K^{2}$ $=k_{j}^{\prime}k^{\prime j}=k_{\nu}\Lambda_{\ j}^{\nu}\Lambda_{\ \mu}^{j}k^{\mu}.$ (39)
As expected, Eq. (37) says that $x^{\prime\prime}$ is also a rest frame. In the frame $x^{\prime\prime},$ the matrix $M^{\prime\prime}\equiv M\left(k^{\prime\prime},u_{0}^{\prime\prime},\omega_{p}\right)$ is simplified to
$M^{\prime\prime}=\left(\begin{array}[]{ccccccccc}\Omega^{2}-K^{2}&0&0&0&\omega_{p}^{2}&0&0&0&\omega_{p}^{2}\\ 0&\Omega^{2}-K^{2}&0&0&0&\omega_{p}^{2}&0&0&0\\ 0&0&\Omega^{2}-K^{2}&0&0&0&\omega_{p}^{2}&0&0\\ 0&0&0&\Omega^{2}-K^{2}&0&0&0&\omega_{p}^{2}&0\\ 0&0&0&0&-\Omega&0&0&0&0\\ K&-\Omega&0&0&0&-\Omega&0&0&0\\ 0&0&-\Omega&0&0&0&-\Omega&0&0\\ 0&0&0&-\Omega&0&0&0&-\Omega&0\\ 0&0&0&0&-\Omega&K&0&0&-\Omega\end{array}\right).$ (40)
It is straightforward to calculate the dispersion relation in the frame $x^{\prime\prime}$ from $M^{\prime\prime}$,
$D^{\prime\prime}\equiv\det\left(M^{\prime\prime}\right)=-\Omega^{3}\left(\Omega^{2}-K^{2}\right)^{2}\left(\Omega^{2}-\omega_{p}^{2}\right)\left(\Omega^{2}-K^{2}-\omega_{p}^{2}\right)^{2}.$ (41)
Detailed calculations show that the spectrum of $M^{\prime\prime}$ consists of the following 5 classes.
  1. 1.
    Photons and anti-photons. For the eigenmodes at
    $\Omega=\sqrt{\omega_{p}^{2}+K^{2}}\ \ \text{and}\ \ \Omega=-\sqrt{\omega_{p}^{2}+K^{2}},$ (42)
    the corresponding eigenvector space for the modes is
    $\left\{\left(0,0,0,-1\right),\left(0,0,0,1\right),0\right\}\oplus\left\{\left(0,0,-1,0\right),\left(0,0,1,0\right),0\right\},$ (43)
    which is 2-dimensional. There is no density perturbation, and in this co-moving frame, the modes are transverse, i.e.,
    $\delta\boldsymbol{E}^{\prime\prime}\cdot\boldsymbol{k}^{\prime\prime}=\delta E_{j}^{\prime\prime}k^{\prime\prime j}=0.$ (44)
    Because of these polarization characteristics, we identify the mode at $\Omega=\sqrt{\omega_{p}^{2}+K^{2}}$ as the photon and the mode at $\Omega=-\sqrt{\omega_{p}^{2}+K^{2}}$ as the anti-photon.
  2. 2.
    Plasmons and anti-plasmons. There are two modes at
    $\Omega=\omega_{p}\ \ \text{and}\ \ \Omega=-\omega_{p}.$ (45)
    The eigenvector space for the $\Omega=\omega_{p}$ mode is
    $\left\{\frac{\omega_{p}^{2}}{K^{2}-\omega_{p}^{2}}\left(1,\frac{\omega_{p}}{K},0,0\right),\left(0,\frac{\omega_{p}}{K},0,0\right),1\right\},$ (46)
    and that for the $\Omega=-\omega_{p}$ mode is
    $\left\{\frac{\omega_{p}^{2}}{K^{2}-\omega_{p}^{2}}\left(1,-\frac{\omega_{p}}{K},0,0\right),\left(0,-\frac{\omega_{p}}{K},0,0\right),1\right\}.$ (47)
    In this co-moving frame, both modes are electrostatic and longitudinal, i.e.,
    $\delta\boldsymbol{B}^{\prime\prime}$ $=0,$ (48)
    $\boldsymbol{k}^{\prime\prime}\times\delta\boldsymbol{E}^{\prime\prime}=\epsilon_{ijl}k^{\prime\prime j}\delta E^{\prime\prime l}$ $=0.$ (49)
    We identify the first mode as the plasmon and the other as the anti-plasmon.
  3. 3.
    Current-plasmons. Two of the eigenvectors at $\Omega=0$ are magnetostatic perturbations generated by plasma current perturbation. Their eigenvector space is
    $\left\{\left(0,0,0,\frac{\omega_{p}^{2}}{K^{2}}\right),\left(0,0,0,1\right),0\right\}\oplus\left\{\left(0,0,\frac{\omega_{p}^{2}}{K^{2}},0\right),\left(0,0,1,0\right),0\right\}.$ (50)
    The polarization directions of the modes are similar to that of photons, but the phase between the current and vector potential is shifted by $\pi.$ The relative amplitude between the magnetic field and normalized flow velocity is proportional to the plasma density, akin to ratio between electrical field and normalized density in the case of plasmons. In this sense, this 2D mode is a hybrid between photons and plasmons. We will call it current-plasmon. It was not identified in previous studies, probably because it has a zero frequency in the rest frames $x^{\prime}$ and $x^{\prime\prime}.$ But in the lab frame $x$, the mode is electromagnetic, oblique, and propagating (see below).
  4. 4.
    Gauge modes. The eigenmodes at $\Omega=\pm K$ have the same dispersion relations as the photon and anti-photon in vacuum. The eigenvectors for $\Omega=K$ and $\Omega=-K$ are, respectively,
[…]
    These modes carry no electric, magnetic, velocity, or density perturbations. They are the gauge modes introduced by the unfixed gauge freedom, and will be regarded as trivial modes that have no observable effects in the classical regime.
  5. 5.
    Off-shell virtual mode. Since $\Omega=0$ is a triple root of the $D^{\prime\prime}=0,$ in addition to the 2D current-plasmon identified above, the system admits another eigenvector at $\Omega=0$, which is
    $\left\{\left(0,0,0,0\right),\left(-1,0,0,0\right),1\right\}.$ (52)
    However, it does not satisfy the on shell condition (21). It is not a valid physical mode in the classical regime and will be discarded.
The classification of the electromagnetic excitations described above resembles Wigner’s little group method for the classification of massive particles [12, 13, 14, 15], except that the base manifold in the present study is the product of the space of equilibrium flow $u_{0}$ and the space of 4-wavevector $k$. The Lorentz transformation applies to $u_{0}$ and $k$ simultaneously.
The three modes identified, i.e., photons, plasmons, and current plasmons, were classified according to their properties in the specially chosen co-moving frame $x^{\prime\prime}$. But the labels for the classification, i.e., transversality, longitudinality, electrostaticity, and magnetostaticity, are not covariant. In the lab frame $x$, the polarizations and dispersion relations are expected to be very different. Because the theory is manifestly covariant, the dispersion relations and eigenvectors in the lab frame $x$ can be easily obtained via the Lorentz transformation,
$\left(\Omega,K,0,0\right)^{T}$ $=k^{\prime\prime}=\Lambda^{\prime}\Lambda k,$ (53)
$\psi$ $=\Lambda^{-1}\Lambda^{\prime-1}\psi^{\prime\prime}.$ (54)
Specifically, pushing forward the 4-wavevector $k$ in the lab frame $x$, i.e., expressing the dispersion relation in the frame $x^{\prime\prime}$ in terms of $k$ via Eq. (53), gives the dispersion relation in the lab frame $x$. The pull-back of $\psi^{\prime\prime}$ in the fame $x^{\prime\prime}$ according to Eq. (54) is the eigenvector $\psi$ in the lab frame $x$. We list below the properties of photons, plasmons, and current-plasmons in the lab frame $x$.
  1. 1.
    Photons and anti-photons in the lab frame. Simple calculation according to the push-forward algorithm given above shows that the dispersion relations for photons and anti-photons in the lab frame $x$ are, respectively,
    $\omega=\sqrt{\omega_{p}^{2}+k_{j}k^{j}}\ \ \text{and}\ \ \omega=-\sqrt{\omega_{p}^{2}+k_{j}k^{j}},$ (55)
    which is expected because $\omega^{2}-k_{j}k^{j}$ is a Lorentz invariant scalar. The dispersion relations are identical to their counterparts, Eq. (42), in the rest frame $x^{\prime\prime}.$ A direct consequence of this result is that the cutoff frequency for photons in a relativistic plasma does not depend on how fast the plasma is moving in the lab frame $x$. The manifest covariance of the dispersion relation demonstrates that relativistic transparency [1, 2, 7, 8, 3, 4, 5, 9, 6], despite being widely studied, does not exist in plasmas. However, the polarization properties of the eigenmodes change under the Lorentz transformations. In the frame $x^{\prime},$ photons are still transverse because $\Lambda^{\prime-1}\in\text{SO}(3),$ which rotates both $\delta\boldsymbol{E}^{\prime\prime}$ and $\boldsymbol{k}^{\prime\prime}$as 3D vectors and preserves the transversality. On the other hand, $\Lambda^{-1}$ is a boost, and it transforms $\boldsymbol{k}^{\prime}$ as the spatial components of the contravariant 4-vector $k^{\prime}$ and $\delta\boldsymbol{E}^{\prime}$ as the temporal components of the electromagnetic tensor $F$. The transversality is broken by a general boost $\Lambda^{-1}$ unless $\omega_{p}$ vanishes. To see this, let’s take the example of
    $\psi^{\prime}$ =$=\left\{\left(0,0,0,-1\right),\left(0,0,0,1\right),0\right\},$ (56)
    $k^{\prime}$ $=(\sqrt{\omega_{p}^{2}+K^{2}},K,0,0),$ (57)
    which is a transverse photon in the rest frame $x^{\prime}$ propagating along the 1st-direction with electrical field polarized in the $3$rd-direction. Assume the plasma flow in the lab frame $x$ is in the 3rd-direction, i.e.,
    $\Lambda^{-1}=\Lambda\left(0,0,-\beta\right),$ (58)
    then in the lab frame $x$, we find
    $\cos\theta_{1}\equiv\frac{\left|\delta\boldsymbol{E}\cdot\boldsymbol{k}\right|}{\left|\delta\boldsymbol{E}\right|\left|\boldsymbol{k}\right|}=\frac{\gamma\left|\beta\right|}{\sqrt{\left(\frac{\omega^{2}}{\omega_{p}^{2}}-\beta^{2}\gamma^{2}\right)\left(\frac{\omega^{2}}{\omega_{p}^{2}}-1\right)}}\leq 1,$ (59)
    where $\theta_{1}$ is the pitch angle between $\delta\boldsymbol{E}$ and $\boldsymbol{k}$. This expression shows that photons carry an electrical field in the wavevector direction unless the plasma disappears ($\omega_{p}=0$) or is motionless $(\beta=0)$. The maximum value of $\cos\theta_{1}$ is reached at $\omega/\omega_{p}=\gamma$, which is the smallest value of $\omega/\omega_{p}$ under the boost in the 3rd-direction. In general, photons and anti-photons in plasmas are neither transverse nor longitudinal; they are oblique.
  2. 2.
    Plasmons and anti-plasmons in the lab frame. By the same push-forward algorithm above, the dispersion relations for plasmons and anti-plasmons in the lab frame $x$ are found to be
    $\omega=k^{j}\beta_{j}+\frac{\omega_{p}}{\gamma}\ \ \text{and}\ \ \omega=k^{j}\beta_{j}-\frac{\omega_{p}}{\gamma},$ (60)
    which can be written in a manifestly covariant manner as
    $k^{\mu}u_{0\mu}=-\omega_{p}\ \ \text{and}\ \ k^{\mu}u_{0\mu}=\omega_{p}.$ (61)
    Importantly, in the lab frame $x$, plasmons are electromagnetic and oblique as photons. For example, take the electrostatic longitudinal plasmon in the rest frame $x^{\prime}$ with
    $\psi^{\prime}$ =$=\left\{\frac{\omega_{p}^{2}}{K^{2}-\omega_{p}^{2}}\left(1,\frac{\omega_{p}}{K},0,0\right),\left(0,\frac{\omega_{p}}{K},0,0\right),1\right\},$ (62)
    $k^{\prime}$ $=(\omega_{p},K,0,0).$ (63)
    Assume again that the plasma flow is in the 3rd-direction in the lab frame $x$, with the Lorentz matrix given by Eq. (58). In the lab frame $x,$ the plasmon has a non-vanishing magnetic component in the 2nd-direction,
    $\delta\boldsymbol{B}=(0,-i\beta\gamma,0)^{T},$
    and the cosine of the pitch angle $\theta_{2}$ between $\delta\boldsymbol{E}$ and $\boldsymbol{k}$ is
    $\cos\theta_{2}\equiv\frac{\left|\delta\boldsymbol{E}\cdot\boldsymbol{k}\right|}{\left|\delta\boldsymbol{E}\right|\left|\boldsymbol{k}\right|}=\sqrt{1-\beta^{2}\gamma^{2}\omega_{p}^{2}/\boldsymbol{k}^{2}},$ (64)
    which is neither longitudinal nor transverse when $\omega_{p}\neq 0.$
  3. 3.
    Current-plasmons in the lab frame. Finally, the dispersion relations for the current plasmons in the lab frame $x$ is
    $\omega-\beta_{j}k^{j}=0,$ (65)
    whose manifestly covariant form is
    $k^{\mu}u_{0\mu}=0.$ (66)
    Note that the current-plasmon is not simply the magnetostatic current mode in the rest frame Doppler-shifted into the lab frame $x$. In the lab frame $x$, the mode carries a time-dependent electrical component that does not exist in the rest frame. To see that this is an oblique, propagating electromagnetic excitation, we take the example of
    $\psi^{\prime}$ =$=\left\{\left(0,0,0,\frac{\omega_{p}^{2}}{K^{2}}\right),\left(0,0,0,1\right),0\right\},$ (67)
    $k^{\prime}$ $=(0,K,0,0),$ (68)
    which is a static current-plasmon in the rest frame $x^{\prime}$ with wavenumber in the 1st-direction and magnetically polarized in the $2$nd-direction. Assuming the plasma flow in the lab frame $x$ is in the 1st- and 3rd-directions, i.e.,
    $\Lambda^{-1}=\Lambda\left(-\beta_{1},0,-\beta_{3}\right),$ (69)
    then in the lab frame $x$, we find
    $\delta\boldsymbol{E}=\frac{\omega_{p}^{2}}{K\gamma}(-i\beta_{3},0,\beta_{1})^{T}.$ (70)
    The cosine of the pitch angle $\theta_{3}$ between $\delta\boldsymbol{E}$ and $\boldsymbol{k}$ is
    $\cos\theta_{3}\equiv\frac{\left|\delta\boldsymbol{E}\cdot\boldsymbol{k}\right|}{\left|\delta\boldsymbol{E}\right|\left|\boldsymbol{k}\right|}=\sqrt{\frac{\beta_{3}^{2}}{\left(\gamma^{2}-1\right)\left(1-\beta_{3}^{2}\right)}},$ (71)
    which is neither longitudinal nor transverse in general.
As a potential application of the theory developed, the polarization angles of the oblique photons, plasmons, and current-plasmons given by Eqs. (59), (64) and (71) can be explored as diagnostic measures of density and/or flow velocity of relativistic plasmas.
We now focus on the photons in plasmas. For this purpose, it is necessary to consider plasmons as well because photons and plasmons resonate at $(\omega,\boldsymbol{k})=(\gamma\omega_{p},\gamma\boldsymbol{\beta}\omega_{p}).$ On the other hand, photons do not resonate with current-plasmons, which can thus be neglected in this context. We just proved that both photons and plasmons are oblique and electromagnetic in general. How do we tell them apart then? Is there a Lorentz invariant label that can distinguish photons from plasmons? It turns out that covariant compressibility defined by
$\chi\equiv\delta u_{\mu}k^{\mu}$ (72)
is such a label. From the eigenvector spaces listed above, we can calculate that for photons $\chi=0$ and for plasmons $\chi=\omega_{p}.$ In this sense, plasmons are covariantly compressible electromagnetic excitations, and photons (and current-plasmons) are covariantly incompressible ones. This is a more precise and more general description than the conventional picture of plasmons being electrostatic and longitudinal while photons are electromagnetic and transverse.
Figure 1: Photon topology in vacuum (a) and in a relativistic plasma with $\beta=0.6$ (b). The dispersion relations of photons, anti-photons, plasmons, and anti-plasmons are plotted in blue, orange, green, and red, respectively. The momentum space is shown as 2D. Plotted $k_{1}$, $k_{2}$, and $\omega$ are normalized by $\omega_{p}$. The photon topology in vacuum (a) is singular at the $(\omega,\boldsymbol{k})=(0,\boldsymbol{0})$ Dirac point of the light cone, which is the fundamental reason that photons do not admit a spin angular momentum operator [15]. Plasmas remove the photon’s Dirac point by giving it an effective mass of $\omega_{p}$, but create a tilted Dirac-Weyl point at $(\omega,\boldsymbol{k})=(\gamma\omega_{p},\gamma\beta\omega_{p},0)$ by reviving the virtual longitudinal photon (b).
One of the motivations of the present study is to explore the photon and plasmon topology, which is intimately related to topological waves in continuous media [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], spin angular moment of photons or the non-existence thereof [31, 32, 33, 34, 35, 36, 37, 15], and the non-Hermitian or PT-symmetric classical physics [38, 39, 40, 41, 42, 43, 44, 45]. In particular, the comparison between photon topology in vacuum and in plasmas is revealing. It has now been established that the angular momentum for vacuum photons can’t be split into orbital angular momentum and spin angular momentum parts [32, 34, 35, 36, 37]. Unlike massive particles, photons in vacuum do not admit a spin angular momentum operator. This fact is fundamentally attributed to the topological singularity at the $(\omega,\boldsymbol{k})=(0,\boldsymbol{0})$ Dirac point of the light cone [15], see Fig. 1(a). Is the photon topology nontrivial in plasmas as in vacuum? The dispersion relations of photons in plasma is shown in Fig. 1(b) along with those of plasmons. Observe that the $(\omega,\boldsymbol{k})=(0,\boldsymbol{0})$ Dirac point for photons in vacuum does exist in plasmas anymore because of the gap of $2\omega_{p}$ between photons and anti-photons, which is evident from the dispersion relation given by Eq. (55). This can be viewed as photons gaining an effective mass of $\omega_{p}$ from the plasma. However, a new exceptional point at $(\omega,\boldsymbol{k})=(\gamma\omega_{p},\gamma\beta\omega_{p},0)$ emerges due to the interaction between photons and plasmons. At this exceptional point, the 2D photon eigenmode bundle and the 1D plasma eigenmode bundle are not distinguishable. For this reason, it should be called a Dirac-Weyl point. It should be removed from the base manifold of the 2D photon eigenmode bundle, which renders the base manifold of photon eigenmode bundle non-contractible, akin to the situation of photons in vacuum [15, 25], and the associated vector bundle topology is expected to be nontrivial and interesting. But the topology will be different from the photon topology in vacuum [15], because the exceptional point at $(\omega,\boldsymbol{k})=(\gamma\omega_{p},\gamma\beta\omega_{p},0)$ is a tilted Dirac-Weyl point [25] instead of an upright Dirac point. Plasmas remove the photon’s upright Dirac point by giving it effective mass, but create a tilted Dirac-Weyl point by reviving the virtual longitudinal photon. These topics and the implications on the angular momentum of photons in plasmas will be reported in future publications.
Acknowledgements.
This research was supported by the U.S. Department of Energy (DE-AC02-09CH11466).